HUBBLE SPACE TELESCOPE DAILY REPORT #1844 PERIOD COVERED: 0000Z (UTC) 03/27/97 - 0000Z (UTC) 03/28/97 Daily Status Report as of 087/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6764 (Detection and Mass Determination of Low Mass = Companions) The FGSS were used in both the TRANS and POS modes to make = astrometric measurements of nearby, dwarf, M class star GL791 in order to = measure the mass of its possible low mass companion. The intent is to = improve the mass-luminosity relation for the cool end of the main = sequence. The observations were executed as scheduled, and no anomalies = were noted. 1.2 Completed WF/PC-2 7016 (WF/PC-2 SM97 UV Monitor And Cool Down = Procedure) This operation was used as a post-SM97 cooldown of the WF/PC-2 CCD = to -88C with frequent monitoring of contamination via standard star = GRW+70D582 observations in filter F170W. An internal flat and a dark = calibration image were also taken. The observations were completed as = planed, and no problems were reported 1.3 Completed Two Sets of WF/PC-2 6359 (Core Properties of the Bulges = of Spiral Galaxies) The WF/PC-2 was used to make snapshot survey observations of = spiral galaxy NGC-3054 in order to study its central bulges. The aim is = to investigate the core properties, correlations with global galactic = properties, and the frequency of occurrence of multiple and/or = miscentered nuclei. The observations were completed as scheduled, and no = anomalies were noted. 1.4 Completed NIC/1 7135 (NICMOS Intermediate Focus / Alignment) The NICMOS Camera 1 was used to make observations of NGC-3603. = The purpose of this activity is to perform a series of iterative = adjustments of the NICMOS pupil alignment mechanism to establish the = initial optical alignment in focus and tilt for all three cameras. The = observations were completed as schedule, and no anomalies were reported. 1.5 Completed WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected = AGNs) The WF/PC-2 was used to make high resolution observations of the = known, low, red-shifted, x-ray active galactic nuclei (AGN) MS0624.4-5. = These data will allow a study of the optical properties of an unbiased = sample of AGN host galaxies. The observations were completed as = scheduled, and no problems were reported. 1.6 Completed NIC/1 7042 (NICMOS Fine Optical Alignment) The Near-Infrared Camera #1 was used to make observations of = NGC-3603 in order to obtain fine optical alignment calibration = measurements. This is to perform a series of iterative adjustments of = the NICMOS pupil alignment mechanism to establish the initial optical = alignment in focus and tilt. The observations were executed as = scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 7019 (WF/PC-2 SMOV Flat Field Check) The WF/PC-2 was used to obtain a sequences of Earth streak and = internal flats to check for changes at the 1% level. The observations = were completed as scheduled, and no problems were reported 1.8 Completed WF/PC-2 6352 (Dwarf Elliptical Galaxy Snapshots) The WF/PC-2 was used to survey VCC991, from a sample of 30 dwarf = elliptical galaxies. The observations were completed as planned, and no = anomalies were reported. 1.9 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy IR19254-72. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.10 Completed WF/PC-2 7029 (WF/PC-2 Cycle 6/SMOV Lyman Alpha = Check-Part 2) The WF/PC-2 was used to make observations of the WD0310-688 in = order to obtain far-ultraviolet (Lyman Alpha) measures of the system's = throughput. This is to quantify any contamination of the WF/PC-2 = pick-off mirror caused by SMOV activities. The observations were = completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: - HSTAR 5997 Battery 1 recharger ratio OOL low @87/04:53z. Under review = by system engieering. - HSTAR 5998 NICMOS filter wheel retry message@87:04:10z. Under = investigation. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of planned SMOV activities. /DMH